How many grams of H2 are needed to produce 11.98 g of NH3?
1
Expert's answer
2016-09-16T09:01:15-0400
Production of the NH3 from hydrogen and nitrogen:
3H2 + N2 → 2NH3
3 mol 1 mol 2 mol
The number of moles NH3:
n(NH3) = m(NH3) / M(NH3) = 11.98 g / 17 g/mol = 0.704 mol
The number of moles H2:
n(H2) = (n(NH3) /2) × 3 = ( 0.704 mol / 2 ) × 3 = 1.06 mol
m (H2) = n(H2) × M (H2) = 1.06 mol × 2 g/mol = 2.11 g
Answer: 2.11 g